OpenStudy (anonymous):

7x-2y=-1 what is the slope of a line parallel to this line? what is the slop perpendicular to this line?

OpenStudy (anonymous):

idk but i think the slope is y=7/2x+1/2

OpenStudy (anonymous):

Thats a profile pic of my bro. He is actually going to a modeling agency tomorrow

OpenStudy (anonymous):

omgfff i didnt click it lol sorrry

OpenStudy (anonymous):

y = mx + b 7x-2y=-2 -2y = -7x-2 y = (7/2)x + 1 Since the slope of the line is 7/2, a line perpendicular to this will have a "negative reciprocal" slope. Let m = slope of perpendicular line then.m(7/2) = -1 m = -2/7 . And, I need to know the slope of a line parallel to the line 7x-2y=-2 Since we found the slope of: 7x-2y=-2 to be 7/2 a line parallel to this will have the SAME slope therefore, slope of a parallel line is 7/2